You need to estimate the value of Laputa Aviation. You have the following forecasts (in millions of dollars) of its profits and of its future investments in new plant and working capital Year Earnings before interest, taxes, depreciation, and amortization (EBITDA) Depreciation Pretax profit Tax at 40% Investment 78 38 40 16 14 98 48 50 20 17 113 118 58 60 24 53 60 24 20 From year 5 onward, EBITDA, depreciation, and investment are expected to remain unchanged at year-4 levels. Laputa is financed 40% by equity and 60% by debt, its cost of equity is 13%, its debt yields 9%, and it pays corporate tax at 40% a. Estimate the companys total value. (Do not round intermediate calculations. Enter your answer in millions rounded to the nearest whole dollar amount.) Total value million b. What is the value of Laputas equity? (Do not round intermediate calculations. Enter your answer in millions rounded to 2 decimal places.) Laputas equity million
Answer #1

To estimate the value of the company, present value of the future free cash flows are found at the WACC. Calculation of Free43 Calculation of value of the company: All amount in million $ Year Free Cash Flow 47 1 2 4 $48.00 $61.00 $69.00 $72.00 $72.Formula sheet

A B C D E F G H I J K L
2
3
4 To estimate the value of the company, present value of the future free cash flows are found at the WACC.
5
6 Calculation of Free Cash Flows:
7
8 First free cash flow needs to be calculated using the following formula:
9 Free Cash Flow = Operating Cash Flow - Capex - Change in working capital
10 Operating Cash Flow = EBIT*(1-T)+Depreciation
11
12 Tax Rate 0.4
13
14 Amount in million$
15 Year 0 1 2 3 4 5
16 EBITDA 78 98 113 118 =H16
17 Depreciation 38 48 53 58 =H17
18 EBIT =E16-E17 =F16-F17 =G16-G17 =H16-H17 =I16-I17
19 Tax expense =-E18*$D$12 =-F18*$D$12 =-G18*$D$12 =-H18*$D$12 =-I18*$D$12
20 EBIT*(1-T) =SUM(E18:E19) =SUM(F18:F19) =SUM(G18:G19) =SUM(H18:H19) =SUM(I18:I19)
21 Add: Depreciation =E17 =F17 =G17 =H17 =I17
22 Operating Cash Flow = EBIT*(1-T)+depreciation =E20+E21 =F20+F21 =G20+G21 =H20+H21 =I20+I21
23 Investment in working capital 0 0 0 0 0
24 Capex 14 17 20 22 =H24
25 Free Cash Flow =E22-E23-E24 =F22-F23-F24 =G22-G23-G24 =H22-H23-H24 =I22-I23-I24
26
27
28 Calculation of WACC:
29 Formula for WACC is given as:
30 WACC = r(E) × w(E) + r(D) × (1 – t) × w(D)
31 Where, r(E) and r(D) are cost of equity and cost of debt, w(E) is weight of equity and W(D) is weight of debt and t is the tax rate
32
33 Tax rate =D12
34 Source of capital Weight(w) Cost(c)
35 Debt 0.6 0.09
36 Equity =1-D35 0.13
37
38 WACC = r(E) × w(E) + r(D) × (1 – t) × w(D)
39 =E36*D36+E35*(1-D33)*D35 =E36*D36+E35*(1-D33)*D35
40
41 Hence WACC is =D39
42
43
44 Calculation of value of the company:
45
46 All amount in million $
47 Year 0 1 2 3 4 5
48 Free Cash Flow =E25 =F25 =G25 =H25 =I25
49
50 Since cash flow after year 4 is constant, value of terminal cash flow at the end of year 4 can be calculated as follows:
51 Terminal cash flow in year 4 =FCF5/w
52 =I48/D41 =I48/D41
53
54 All amount in million $
55 Year 0 1 2 3 4
56 Free cash flow =E48 =F48 =G48 =H48
57 Horizon value =D52
58 WACC =D41
59 Present Value Factor (P/F,i,n) =1/(1+$D58)^E55 =1/(1+$D58)^F55 =1/(1+$D58)^G55 =1/(1+$D58)^H55
60 Present value of cash flows =E56*E59 =F56*F59 =G56*G59 =(H56+H57)*H59
61 Present value of cash flows =SUM(E60:H60) =SUM(E60:H60)
62
63 Value of Firm =D61 million
64 Value of Debt =D63*D35 million
65 Value of Equity =D63-D64 million
66
67 Hence Value of the firm is =D63 million
68
69 Value of Equity =D65 million
70